OpenStudy (anonymous):

The unit price of x units of a certain product is given by P(x)=(450/x+3)-3. What is the maximum possible revenue when selling x units?

OpenStudy (anonymous):

Is this a calculus problem?

OpenStudy (anonymous):

\[ P(x) = \frac{450}{x+3}-3 \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it a calculus class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Find the derivative first.

OpenStudy (anonymous):

I found it and got about $332, but I'm not sure.

OpenStudy (anonymous):

Did you get the derivative?

OpenStudy (anonymous):

Yes, -450/(x+3)^2

OpenStudy (kropot72):

The maximum revenue from selling x units is required. Let the revenue be R. To find the revenue from selling x units, we need to multiply the unit price by x, giving: \[\large R(x)=\frac{450x}{x+3}-3x\ .........(1)\] Are you with me so far?

OpenStudy (kropot72):

@kperez128 Are you there?

OpenStudy (anonymous):

yes, I did that

OpenStudy (kropot72):

Now you need find dR/dx. If you have done that what was your result?

OpenStudy (anonymous):

\[\frac{ 450x}{x+3^{2} }+\frac{ 450 }{ x+3}-3\]

OpenStudy (kropot72):

My result of the differentiation of (1) with respect to x is as follows: \[\large \frac{dR}{dx}=\frac{450(x+3)-450x}{(x+3)^{2}}-3\ .........(2)\]

OpenStudy (kropot72):

@kperez128 Are you checking your differentiation?